Consider a deposit of $1,300 made today in an investment account. The account yields an annual return of 3.90% for the next 3 years, then returns 6.90% annually for 9 years after that, and finally yields 2.30% annually for another 8 years, What will be the balance of this account (rounded to the nearest dollar) at the end of this 20 year period? Assume no additional withdrawals or deposits during this period.
Answer #1

We use the formula:  
A=P(1+r/100)^n
where   
A=future value
P=present value  
r=rate of interest
n=time period.

A=1300*(1.039)^3*(1.069)^9*(1.023)^8

=1300*2.45273743

=$3189(Approx)